Solve
a) log(base 2)(x-5) + log(base 2) (x-2) = 2
Can you please help me out? Thanks so much in advance:)

Answer by nerdybill(7384)

log(base 2)(x-5) + log(base 2) (x-2) = 2
.
log(2,x-5) + log(2,x-2) = 2
log(2,(x-5)(x-2)) = 2
(x-5)(x-2) = 2^2
(x-5)(x-2) = 4
x^2-7x+10 = 4
x^2-7x+6 = 0
(x-6)(x-1) = 0
.
x = {1, 6}
You have to throw out 1 (extraneous) leaving
x = 6